OpenStudy (anonymous):

determine if the equation is a function y2=3-x2

OpenStudy (anonymous):

Is this y^2=3-x^2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y = sqrt(3-x^2) this is a one- to -many relation - this should help

OpenStudy (anonymous):

\[y= \pm \sqrt{3-x^2}\]. For every x, there will be 2 y values, not a function

OpenStudy (anonymous):

ok not a function was the answer i was looking for thank you

OpenStudy (anonymous):

yw